40'x 80' rectangular lot...2 boys mowing one decided that will stop at a uniform distance ..moving using the collapsing pattern....what uniform distance should one stop mowing
Answer by edjones(8007)

40*80=3200 area
3200/2=1600 half area each must mow.
Let x=the uniform distance from the perimeter.
(40-x)(80-x)=1600
3200-120x+x^2=1600
x^2-120x+1600=0
x^2-120x+3600=3600-1600 completing the square.
(x-60)^2=2000
x-60=+-sqrt(400*5)
x=60+20sqrt(5)=104' approx (too large)
x=60-20sqrt(5) ' Answer (15.28' approx)
.
Ed